# Large Language Model Aided QoS Prediction for Service Recommendation

### v1

@@ Abstract
Large language models (LLMs) have seen rapid improvement in the recent years, and are used in a wider range of applications. After being trained on large text corpus, LLMs obtain the capability of extracting rich features from textual data. Such capability is potentially useful for the web service recommendation task, where the web users and services have intrinsic attributes that can be described using natural language sentences and are useful for recommendation. In this paper, we explore the possibility and practicality of using LLMs for web service recommendation. We propose the large language model aided QoS prediction (llmQoS) model, which use LLMs to extract useful information from attributes of web users and services via descriptive sentences. This information is then used in combination with the QoS values of historical interactions of users and services, to predict QoS values for any given user-service pair. Our proposed model is shown to overcome the data sparsity issue for QoS prediction. We show that on the WSDream dataset, llmQoS outperforms comparable baseline models consistently.
@@ I Introduction
The rapid development of cloud computing has led to an explosive growth in the number of cloud services, resulting in many services having identical or similar functionalities [1]. There is a growing need for service providers to provide the most suitable service for users to meet customer demands. Where there are many candidate services for a user request, quality of service (QoS) is a commonly used metric to measure the performance of these services. QoS refers to non-functionality aspects of services [2] such as throughput, response time, and cost. It plays a crucial role in distinguishing similar services and helps users select the most appropriate one. However, in real-world scenarios, evaluating the QoS of a large number of services on the client side is expensive and time-consuming due to the sheer scale. Additionally, most end users lack the expertise to do the evaluation [3]. Therefore, the only feasible option for an end user is to predict the QoS values for an unknown service, and the QoS values are used to help the user find the most suitable service.
@@
The problem of QoS prediction is illustrated in Figure I. For the sets of web users and services exists a QoS matrix, in which each value corresponds to a user-service interaction. The matrix contains some known QoS values from historical user-service interactions, and our goal is to predict unknown QoS values using these known values, thereby recommending more suitable services to users. Many methods have been developed to predict the QoS values for a given user and service. Among them, collaborative filtering is the most widely used technique. It aim to accurately estimate QoS values by learning from the QoS value matrix from historical user-service interactions. However, in practice, the historical QoS value matrix is often highly sparse, as each user usually only interacted with a limited number of cloud services in the past. This data sparsity issue poses difficulty for many QoS prediction methods. Although many methods aim to mitigate this difficulty, it is still challenging to achieve ideal predictive performance.
@@
More recently, large language models (LLMs) have demonstrated remarkable performance in understanding and generating texts. LLMs have large capacity and are trained on vast amounts of data, enabling them to capture complex patterns in natural language with a long context. They have rapidly become powerful tools for various natural language processing tasks. For the QoS prediction problem, each user and service has some specific attributes. For instance, users may have attributes such as its country of origin and the autonomous system it belongs to. Given that LLMs are capable of handling natural language with contextual information, we claim that such capability can be used for QoS prediction task as well. In this paper, we propose to use natural language sentence to describe each user or service using their corresponding attributes. Those attributes contain information such as the proximity and type, which are potentially useful for QoS prediction. We then use state-of-the-art LLMs to extract the descriptive features of users and services from those sentences. This LLMs-based feature helps the QoS prediction model better understand and represent the latent features of users and services. It effectively mitigates the data sparsity issue, as the descriptive LLM features does not rely on historical user-service interactions. It helps the model to provide more reliable and effective service recommendations.
@@
  • We first introduce large language models for the web service recommendation task. We propose large language model aided QoS prediction (llmQoS), a novel approach that combines collaborative filtering and nature language processing, making use of both historical user-service interaction information and rich text feature from attributes of users and services. It can effectively mitigate the data sparsity issue inherent to the QoS prediction problem.
  • We show that llmQoS can predict QoS values accurately under different data sparsity levels. The llmQoS model is robust, and can work with different LLMs and even collaborative filtering network architectures. On the WSDream dataset, it outperforms several existing QoS prediction models consistently.
@@ II-A QoS Prediction for Service Recommendation
Memory-based CF methods first calculate the similarity of users or services, and then predict the QoS values of the target user and service based on historical QoS values and the similarity relationships with their neighbors. Shao et al. [5] first utilize user-based CF methods for QoS prediction. Qi et al. [6] identify “enemies” of the target user to determine their “potential friends”, successfully overcoming the cold-start problem. Meanwhile, item-based methods are widely employed [7, 8]. Chen et al. [9] propose a predictive method that integrates service similarity and geographical location information, effectively alleviating data sparsity and cold-start issues. Additionally, hybrid algorithms that compute both service-to-service and user-to-user similarities [10] are developed. Zheng et al. [11] simultaneously consider user and service similarities for QoS prediction, demonstrating superior predictive performance compared to previous methods.
@@
Deep learning technology has demonstrated remarkable capabilities across various fields of artificial intelligence in recent years. And it is gradually gaining more attention in the field of QoS prediction too. An increasing number of deep learning based techniques for QoS prediction are developed [15, 16]. Zou et al. [17] propose a novel adaptive QoS prediction framework, utilizing a dual-tower deep residual network to extract latent features of users and services and predict QoS values. Wu et al. [18] introduce a reputation-integrated graph convolutional network for robust and accurate QoS prediction. Zhu et al. [19] develop a twin graph network based on graph contrastive learning, effectively addressing cold-start and data sparsity issues.
Most deep learning based QoS prediction models use the user and service ID information through embedding layers. Additionally, some methods [17, 19] utilize more information of the users and services such as their location, IP address, and the autonomous systems they belong to. Such information provides similarity or proximity signal for the model, which are useful for QoS prediction. In our work, we utilize such information through large language models.
@@ II-B Large Language Model
In the field of natural language processing (NLP), large language models (LLMs) have seen rapid development since the introduction of the transformer architecture [20]. They have shown state-of-the-art performance across virtually all NLP tasks, including text classification and summarizing, machine translation, text generation, and conversation systems.
@@ III-A Large Language Model Feature Extraction
RoBERTa and Phi3mini have similar overall architecture. Each LLM layer in them comparably consists of an attention based layer [20] and a set of feed-forward layers. The attention layer calculates the correlation among the feature vectors in the sequence. The feed-forward layers are combination of linear layers, normalization layers, and activation functions.
But RoBERTa and Phi3mini have different number of layers, different dimensionality of the embedding vectors, and different tokenizer. Phi3mini has $3.5\times 10^{9}$ parameters, which is significantly larger than RoBERTa’s $1.2\times 10^{8}$ million parameters. They are also pre-trained on different datasets using different training tasks. RoBERTa is pre-trained to do text classification on total of 160 giga-bytes of text data. If each token uses 5 bytes on average, the model sees $3.2\times 10^{10}$ tokens during training. For text classification, the first vector in the final sequence is used. So we also use the first vector as the sentence-level feature. On the other hand, Phi3mini is pre-trained using next token prediction on a dataset with $3.3\times 10^{12}$ tokens. Next token prediction uses the last vector in the final sequence, so we also use the last vector in the final sequence for Phi3mini. Later we show in experiments that Phi3mini outperforms RoBERTa consistently.
@@ III-B Model Architecture
Please note that our network design is straightforward without complicated architectures such as graph convolution or attention. The QoS prediction network architecture is not the focus of this paper. With straightforward architecture, llmQoS can still achieve state-of-the-art performance as shown by the experiments. This demonstrate the effectiveness of incorporating LLM features for QoS prediction. In § IV-G we further demonstrate that LLM features can be integrated into more complex QoS prediction models and can still increase the performance significantly and consistently.
@@ IV Experiments
Model Response time
@@ IV-C Baseline Models
FIGURE VI: QoS prediction performance in terms of Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of BGCL model with different LLM feature extractors. We compare Phi3mini, RoBERTa, and vanilla BGCL. Similar to Figure V, LLM features are beneficial for QoS prediction, and while larger Phi3mini achieves better QoS prediction performance than RoBERTa.
@@ IV-D Comparison with Baselines
The performance of the proposed llmQoS is compared with the baseline methods in Table I. Phi3mini is used for llmQoS in this experiment. We compare both MAE and RMSE for throughput and response time at different densities. It is evident that llmQoS outperforms all baselines consistently at all densities. For throughput value prediction, llmQoS can reduce the MAE by more than 20%, and RMSE by more than 6% from the best baseline model. For response time value prediction, llmQoS can also reduce MAE for more than 10% from the best baseline method.
The results demonstrate that with the rich features from the descriptive sentences extracted by LLMs, the QoS prediction model can effectively find the similarity and correlation among users and services. Along with the information learned from historical interactions, llmQoS can achieve high QoS prediction accuracy.
@@ IV-F Comparison of Different LLMs
IV-G Adding LLM Feature to BGCL
@@
We keep all other features and network architecture unchanged for BGCL, with the only addition of the LLM feature vectors along side with the embedding vectors of BGCL at the very beginning of the network of BGCL. Similar to § IV-F, we compare BGCL with Phi3mini, BGCL with RoBERTa, and vanilla BGCL. The experimental results, as shown in Figure VI, are similar to the results in Figure V too. We further verify that that integrating LLM features significantly enhances the model’s performance, and larger model achieves more satisfactory results. This demonstrates that LLM features are beneficial to QoS prediction problem, regardless of the network architecture. And we expect this to apply to other QoS prediction models as well.
IV-H Hyper-Parameter Tuning of MLP
(a) Throughput / depth
(b) Response / depth
(c) Throughput / width
(d) Response time / width
FIGURE VII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model different depth and width configurations for the MLP network (§ III-B). Depth refers to the number of layers in the MLP, the main experiments use 3 layers. For different width configurations, we multiply the dimensionalities of each layers used by the main experiments by different width factor for comparison. We calculate and show the relative change of MAE and RMSE of a depth or width configuration from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. Generally, the effect of configuration of MLP has on the QoS prediction performance is insignificant, with changes less than 5% in most cases.
We test the performance of llmQoS using different configurations for the MLP model as described in § III-B. For this experiment we test the density level $D$=5%, and using Phi3mini as the LLM feature extractor. We vary the depth of the MLP by adding or removing layers from it. The QoS prediction performance of different configurations is shown in 7(a) and 7(b). We change the width of the MLP by multiplying a width factor to the dimensionalities of the layers in it. The QoS prediction performance of different configurations is shown in 7(c) and 7(d).
@@ IV-H Hyper-Parameter Tuning of MLP
IV-I Training Parameter Tuning
(a) Throughput / BS
(b) Response time / BS
(c) Throughput / LR
(d) Response time / LR
FIGURE VIII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model trained using different batch sizes (BS) and learning rates (LR). We calculate and show the relative change of MAE and RMSE of a batch size or learning rate from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. The performance is generally insensitive to batch size. It is sensitive to learning rate, with the optimal value around $10^{-4}$.
@@ IV-I Training Parameter Tuning
The QoS performance of models trained using different batch sizes is compared in 8(a) and 8(b). Generally speaking, the impact of batch size is inconsistent and insignificant. Only for throughput, larger batch size of 1024 causes the prediction accuracy to drop significantly. The performance with different learning rates is compared in 8(c) and 8(d). Learning rate has a huge impact on the model’s prediction accuracy. Very small learning rate results in slow convergence, while large learning rate causes unstable training. Both causes low performance. The optimal learning rate is around $10^{-4}$ as used in the main experiments.

### v2

@@ Abstract
Large language models (LLMs) have seen rapid improvement in the recent years, and have been used in a wider range of applications. After being trained on large text corpus, LLMs obtain the capability of extracting rich features from textual data. Such capability is potentially useful for the web service recommendation task, where the web users and services have intrinsic attributes that can be described using natural language sentences and are useful for recommendation. In this paper, we explore the possibility and practicality of using LLMs for web service recommendation. We propose the large language model aided QoS prediction (llmQoS) model, which use LLMs to extract useful information from attributes of web users and services via descriptive sentences. This information is then used in combination with the QoS values of historical interactions of users and services, to predict QoS values for any given user-service pair. On the WSDream dataset, llmQoS is shown to overcome the data sparsity issue inherent to the QoS prediction problem, and outperforms comparable baseline models consistently.
@@ I Introduction
The rapid development of cloud computing has led to an explosive growth in the number of cloud services, resulting in many services having identical or similar functionalities [1]. There is a growing need for service providers to provide the most suitable service for users to meet customer demands. When there exist many candidate services for a user request, quality of service (QoS) is a commonly used metric to measure the performance of these services. QoS refers to non-functionality aspects of services [2] such as throughput, response time, and cost. It plays a crucial role in distinguishing similar services and helps users select the most appropriate one. However, in real-world scenarios, evaluating the QoS of a large number of services on the client side is expensive and time-consuming due to the sheer scale. Additionally, most end users lack the expertise to do the evaluation [3]. Therefore, it is common for an end user is to predict the QoS values for an unknown service, and the QoS values are used to help the user find the most suitable service.
@@
The problem of QoS prediction is illustrated in Figure I. For the sets of web users (u1 $\sim$ u4) and services (s1 $\sim$ s4) exists a QoS matrix, in which each value corresponds to a user-service interaction. The matrix contains some known QoS values from historical user-service interactions, and our goal is to predict unknown QoS values (marked by “?”) using these known values, thereby recommending more suitable services to users. Many methods have been developed to predict the QoS values for a given user and service. Among them, collaborative filtering is the most widely used technique. It aim to accurately estimate unknown QoS values by learning from the QoS value matrix from historical user-service interactions. However, in practice, the historical QoS value matrix is often highly sparse, as each user usually only interacted with a limited number of cloud services in the past. This data sparsity issue poses difficulty for many QoS prediction methods. Although many methods aim to mitigate this difficulty, it is still challenging to achieve ideal predictive performance.
@@
More recently, large language models (LLMs) [4, 5] have demonstrated remarkable performance in understanding and generating texts. LLMs have large capacity and are trained on vast amounts of data, enabling them to capture complex patterns in natural language with a long context. They have rapidly become powerful tools for various natural language processing tasks [6, 7, 8, 9, 10]. For the QoS prediction problem, each user and service has some specific attributes. For instance, users may have attributes such as its country of origin and the autonomous system it belongs to. Given that LLMs are capable of handling natural language with contextual information, we claim that such capability can be used for QoS prediction task as well. In this paper, we propose to use natural language sentence to describe each user or service using their corresponding attributes. Those attributes contain information such as the proximity and type, which are potentially useful for QoS prediction. We then use state-of-the-art LLMs to extract the descriptive features of users and services from those sentences. This LLMs-based feature helps the QoS prediction model better understand and represent the latent features of users and services. It effectively mitigates the data sparsity issue, as the descriptive LLM features provide information that is complementary to the historical user-service interactions. Along with the historical QoS values, it helps the model to provide more reliable and effective service recommendations.
@@
  • To the best of our knowledge, we first introduce large language models for the web service recommendation task. We propose large language model aided QoS prediction (llmQoS), a novel approach that combines collaborative filtering and nature language processing, making use of both historical user-service interaction information and rich text feature from attributes of users and services. It can effectively mitigate the data sparsity issue inherent to the QoS prediction problem.
  • The proposed llmQoS model is robust and versatile. It can work with different LLMs, and can generalize to different collaborative filtering network architectures.
  • We show that llmQoS can predict QoS values accurately under different data sparsity levels. On the WSDream dataset, it outperforms several existing QoS prediction models consistently.
@@ II-A QoS Prediction for Service Recommendation
For memory-based CF methods, the similarity of users or services is first calculated. Then QoS values of the target user and service are predicted based on both the historical QoS values and the similarity relationships with their neighbors. Shao et al. [12] first utilize user-based CF methods for QoS prediction. Qi et al. [13] identify “enemies” of the target user to determine their “potential friends”, successfully overcoming the cold-start problem. Meanwhile, item-based methods are widely employed [14, 15]. Chen et al. [16] propose a predictive method that integrates service similarity and geographical location information, effectively alleviating data sparsity and cold-start issues. Additionally, hybrid algorithms that compute both service-to-service and user-to-user similarities [17] are developed. Zheng et al. [18] simultaneously consider user and service similarities for QoS prediction, demonstrating superior predictive performance compared to previous methods. Hu et al. [19] propose to integrate time information to improved CF for QoS prediction.
@@
Deep learning technology has demonstrated remarkable capabilities across various fields of artificial intelligence in recent years [23, 24, 25, 4, 5]. And an increasing number of deep learning based techniques for QoS prediction are developed [26, 27] too. Zou et al. [28] propose a novel adaptive QoS prediction framework, utilizing a dual-tower deep residual network to extract latent features of users and services and predict QoS values. Wu et al. [29] introduce a reputation-integrated graph convolutional network for robust and accurate QoS prediction. Wei et al. [30] use a transformer based method to learn the graph structure for time-aware service recommendation. Zhu et al. [31] develop a twin graph network based on graph contrastive learning, effectively addressing cold-start and data sparsity issues. Zhou et al. [32] propose the spatial context-aware time series forecasting (SCATSF) framework by utilizing both temporal and spatial contextual information of users and services. Lian et al. [33] facilitate multi-task learning for training QoS prediction model.
Most deep learning based QoS prediction models use the user and service ID information through embedding layers. Additionally, to improve prediction performance, some methods [28, 31] utilize more information of the users and services such as their location, IP address, and the autonomous systems they belong to. Such information provides similarity or proximity signal for the model, which are useful for QoS prediction. In our work, we utilize such information through large language models.
@@ II-B Large Language Model
In the field of natural language processing (NLP), LLMs have seen rapid development since the introduction of the transformer architecture [34]. They have shown state-of-the-art performance across virtually all NLP tasks, including text classification and summarizing, machine translation, text generation, and human-like conversation systems [4, 5].
@@ III-A Large Language Model Feature Extraction
RoBERTa and Phi3mini have similar overall architecture. Each LLM layer in them comparably consists of an attention based layer [34] and a set of feed-forward layers. The attention layer calculates the correlation among the feature vectors in the sequence. The feed-forward layers are combination of linear layers, normalization layers, and activation functions. This common architecture is used by many other LLMs too.
However, RoBERTa and Phi3mini have different number of layers, different dimensionality of the embedding vectors, and different tokenizer. Phi3mini has $3.5\times 10^{9}$ parameters, which is significantly larger than RoBERTa’s $1.2\times 10^{8}$ million parameters. They are also pre-trained on different datasets using different training tasks. RoBERTa is pre-trained to do text classification on total of 160 giga-bytes of text data. If each token uses 5 bytes on average, the model sees $3.2\times 10^{10}$ tokens during training. For text classification, the first vector in the final sequence is used. So we also use the first vector as the sentence-level feature. On the other hand, Phi3mini is pre-trained using next token prediction on a dataset with $3.3\times 10^{12}$ tokens. Next token prediction uses the last vector in the final sequence, so we also use the last vector in the final sequence for Phi3mini. Later we show in experiments that Phi3mini outperforms RoBERTa consistently.
@@ III-B Model Architecture
Please note that our network design is straightforward without complicated architectures such as graph convolution or attention. The QoS prediction network architecture is not the focus of this paper. With straightforward architecture, llmQoS can still achieve state-of-the-art performance as shown by the experiments. This demonstrate the effectiveness of incorporating LLM features for QoS prediction. In § IV-H we further demonstrate that LLM features can be integrated into more complex QoS prediction models and can still increase the performance significantly and consistently.
@@ IV Experiments
This section shows the results and discussion of our extensive experiments with llmQoS. It starts with the details of the dataset, evaluation metrics, implementation, and the baseline methods. Then the performance of llmQoS is compared with the baselines. And multiple ablation study experiments are conducted to verify that the LLM feature is beneficial for QoS prediction, and can be integrated to other network architectures too.
@@
Model Sub-dataset response time
@@ IV-C Baseline Models
FIGURE VI: QoS prediction performance in terms of Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of model with added random feature vectors as input. Please refer to § IV-G for more details. The models with random feature vectors do not outperform the ID only model. This indicates the performance gain obtained by llmQoS is indeed from the useful LLM feature instead of added model parameters.
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
FIGURE VII: QoS prediction performance in terms of Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of BGCL model with different LLM feature extractors. We compare Phi3mini, RoBERTa, and vanilla BGCL. Similar to Figure V, LLM features are beneficial for QoS prediction, and while larger Phi3mini achieves better QoS prediction performance than RoBERTa.
(a) Throughput / depth
(b) Response / depth
(c) Throughput / width
(d) Response time / width
FIGURE VIII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model different depth and width configurations for the MLP network (§ III-B). Depth refers to the number of layers in the MLP. The main experiments use 3 layers. For different width configurations, we multiply the dimensionalities of each layers used by the main experiments by different width factor for comparison. We calculate and show the relative change of MAE and RMSE of a depth or width configuration from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. Generally, the effect of configuration of MLP has on the QoS prediction performance is insignificant, with changes less than 5% in most cases.
(a) Throughput / BS
(b) Response time / BS
(c) Throughput / LR
(d) Response time / LR
FIGURE IX: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model trained using different batch sizes (BS) and learning rates (LR). We calculate and show the relative change of MAE and RMSE of a batch size or learning rate from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. The performance is generally insensitive to batch size. It is sensitive to learning rate, with the optimal value around $10^{-4}$.
@@ IV-D Comparison with Baselines
The performance of the proposed llmQoS is compared with the baseline methods in Table I. Phi3mini is used for llmQoS in this experiment. We compare both MAE and RMSE for throughput and response time at different densities. It is evident that llmQoS outperforms all baselines consistently at all densities. For throughput value prediction, llmQoS can reduce the MAE by more than 20%, and RMSE by more than 6% from the best baseline model. For response time value prediction, llmQoS can also reduce MAE for more than 10% from the best baseline method. The results demonstrate that with the rich features from the descriptive sentences extracted by LLMs, the QoS prediction model can effectively find the similarity and correlation among users and services. Along with the information learned from historical interactions, llmQoS can achieve high QoS prediction accuracy.
@@ IV-F Comparison of Different LLMs
IV-G Ablation Study Using Random Feature
In § IV-F we have shown that LLM feature improves QoS prediction accuracy, and larger Phi3mini is more effective than smaller RoBERTa. However, the prediction model with LLM feature input has more trainable parameters than the model that only uses ID embedding. What is more, since the feature vectors from Phi3mini have 3,072 dimensions compared to RoBERTa’s 768 dimensions, the projection layer for Phi3mini also has more parameters than RoBERTa. So it is plausible that the performance improvement is due to the increased number of parameters resulting in large model capacity.
To verify this, we train a model that still has 3,072-dimensional feature vectors as input. But those feature vectors are randomly initialized for each user and service, so they do not contain any useful information. However, this model has the same number of parameters as the llmQoS model with Phi3mini feature. We compare the QoS prediction performance of this model with both the ID only model as in § IV-F and the llmQoS with Phi3mini in Figure VI. Generally, the model with random feature does not outperform ID only model. On throughput dataset under $D$=5and $D$=15%, the accuracy is marginally higher than the ID only model. But under other settings the performance actually worsens. We also observe that with random feature, the model converges with much less epochs than using Phi3mini feature. This indicates that more trainable parameters lead to overfitting. We verify that the performance boost of llmQoS is from the useful information encoded in the LLM feature, instead of added parameters of the model.
IV-H Adding LLM Feature to BGCL
@@
We keep all other features and network architecture unchanged for BGCL, with the only addition of the LLM feature vectors along side with the embedding vectors of BGCL at the very beginning of the network of BGCL. Similar to § IV-F, we compare BGCL with Phi3mini, BGCL with RoBERTa, and vanilla BGCL. The experimental results, as shown in Figure VII, are similar to the results in Figure V too. We further verify that that integrating LLM features significantly enhances the model’s performance, and larger model achieves more satisfactory results. This demonstrates that LLM features are beneficial to QoS prediction problem, regardless of the network architecture. And we expect this to apply to other QoS prediction models as well.
IV-I Hyper-Parameter Tuning of MLP
We test the performance of llmQoS using different configurations for the MLP model as described in § III-B. For this experiment we test the density level $D$=5%, and using Phi3mini as the LLM feature extractor. We vary the depth of the MLP by adding or removing layers from it. The QoS prediction performance of different configurations is shown in 8(a) and 8(b). We change the width of the MLP by multiplying a width factor to the dimensionalities of the layers in it. The QoS prediction performance of different configurations is shown in 8(c) and 8(d).
@@ IV-I Hyper-Parameter Tuning of MLP
IV-J Training Parameter Tuning
@@
The QoS performance of models trained using different batch sizes is compared in 9(a) and 9(b). Generally speaking, the impact of batch size is inconsistent and insignificant. Only for throughput, larger batch size of 1024 causes the prediction accuracy to drop significantly. The performance with different learning rates is compared in 9(c) and 9(d). Learning rate has a huge impact on the model’s prediction accuracy. Very small learning rate results in slow convergence, while large learning rate causes unstable training. Both causes low performance. The optimal learning rate is around $10^{-4}$ as used in the main experiments.